graph the line that contains the point P and has slope m. find the slope form of the equation of the line.
p=(8,1) ; M=6

Answer by Alan3354(69443)

Same problem, different numbers.
------------
Find an equation for the line with the given properties, express the equation in general form.
slope = -1/3 ; containing the point (1,6)
===========
Use y-y1 = m*(x - x1) where m is the slope and (x1,y1) is the point.